# Complexity of deep computations via topology of function spaces

### v1

@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
    AA has NIP.
The Independence Property was first isolated by Saharon Shelah in model theory as a dividing line between theories whose models are “tame” (corresponding to NIP) and theories whose models are “wild” (corresponding to IP). See [Shelah:1971, Definition 4.1],[Shelah:1990]. We will discuss this dividing line in more detail in the next section.
@@ 1.2. NIP as a universal dividing line between polynomial and exponential complexity
The particular case of the BFT dichotomy (Theorem 1.5) when $A$ consists of $\{0,1\}$-valued (i.e., $\{\text{Yes},\text{No}\}$-valued) strings was discovered independently, around 1971-1972 in many foundational contexts related to polynomial (“tame”) vs exponential (“wild”) complexity: In model theory, by Saharon Shelah [Shelah:1971],[Shelah:1990],in combinatorics, by Norbert Sauer [Sauer:1972], and Shelah [Shelah:1972, Shelah:1990], and in statistical learning, by Vladimir Vapnik and Alexey Chervonenkis [Vapnik-Chervonenkis:1971, Vapnik-Chervonenkis:1974].
@@
In Shelah’s classification program [Shelah:1990], theories without the independence property (called NIP theories, or dependent theories) have a well-behaved, “tame” structure; the number of types over a set of parameters of size $\kappa$ of such a theory is of polynomially or similar “slow” growth on $\kappa$. In contrast, Theories with the Independence Property (called IP theories) are considered “intractable” or “wild”. A theory with the Independence Property produces the maximum possible number of types over a set of parameters; for a set of parameters of cardinality $\kappa$, the theory will have $2^{2^{\kappa}}$-many distinct types.
@@
Sauer [Sauer:1972] and Shelah [Shelah:1972] proved the following: If $\mathscr{F}=\{S_{0},S_{1},\dots\}$ is a family of subsets of some infinite set $S$, then either for every $n\in\mathbb{N}$, there is either a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A):i\in\mathbb{N}\}|=2^{n}$ (yielding exponential complexity), or there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
$|\{S_{i}\cap A):i\in\mathbb{N}\}|\leq\sum_{i=0}^{N-1}\binom{|A|}{i}\approx O(|A|^{N})$
(yielding polynomial complexity). This answered a question of Erdős.
@@
Readers familiar with statistical learning may recognize the Sauer-Shelah lemma as the dichotomy discovered and proved slightly earlier (1971) by Vapknis and Chervonenkis [Vapnik-Chervonenkis:1971, Vapnik-Chervonenkis:1974] to address the problem of uniform convergence in statistics. The least integer $N$ given by the preceding paragraph, when it exists, is called the VC-dimension of $\mathscr{F}$. This is a core concept in machine learning. If such an integer $N$ does not exist, we say that the VC-dimension of $\mathscr{F}$ is infinite. The lemma provides upper bounds on the number of data points (sample size $m$) needed to learn a concept class with VC dimension $d\in\mathbb{N}$ by showing this number grows polynomially with $m$ and $d$ (namely, $\sum_{i=0}^{d}{m\choose i}\approx O(m^{d})$), not exponentially. The Fundamental Theorem of Statistical Learning states that a hypothesis class is PAC-learnable (PAC stands for “Probably Approximately Correct”) if and only if its VC dimension is finite.
@@ 1.3. Rosenthal compacta
The comprehensiveness of Theorem 1.5, attested by the examples outlined in the preceding section, led to the following definition (isolated by Gilles Godefroy [Godefroy:1980]):
@@
A Rosenthal compactum is a compact Hausdorff topological space $K$ that can be topologically embedded as a compact subset into the space of all functions of the first Baire class on some Polish space $X$, equipped with the topology of pointwise convergence.
Rosenthal compacta are characterized by significant topological and dynamical tameness properties. They play an important role in functional analysis, measure theory, dynamical systems, descriptive set theory, and model theory. In this paper, we introduce their applicability in deep computation. For this, we shall first focus on countable languages, which is the theme of the next subsection.
@@
Our goal now is to characterize relatively compact subsets of $B_{1}(X,Y)$ for the particular case when $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable. Given $P\in\mathcal{P}$ we denote the projection map onto the $P$-coordinate by $\pi_{P}:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}$. From a high-level topological interpretation, the next lemma states that, in this context, the spaces $\mathbb{R}$ and $\mathbb{R}^{\mathcal{P}}$ are really not that different, and that if we understand the Baire class 1 functions of one space, then we also understand the functions of both.
Lemma 1.10.
Let $X$ be a Polish space and $\mathcal{P}$ be a countable set. Then, $f\in B_{1}(X,\mathbb{R}^{\mathcal{P}})$ if and only if $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$.
@@
Only one implication needs a proof. Suppose that $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$. Let $V$ be a basic open subset of $\mathbb{R}^{\mathcal{P}}$. That is, there exists a finite $\mathcal{P}^{\prime}\subseteq\mathcal{P}$ such that $V=\bigcap_{P\in\mathcal{P}^{\prime}}\pi_{P}^{-1}[U_{P}]$ where $U_{P}$ is open in $\mathbb{R}$. Then,
$f^{-1}[V]=\bigcap_{P\in\mathcal{P}^{\prime}}(\pi_{P}\circ f)^{-1}[U_{P}]$
is an $F_{\sigma}$ set. Since $\mathcal{P}$ is countable, $\mathbb{R}^{\mathcal{P}}$ is second countable so every open set $U$ in $\mathbb{R}^{\mathcal{P}}$ is a countable union of basic open sets. Hence, $f^{-1}[U]$ is $F_{\sigma}$. ∎
Below we consider $\mathcal{P}$ with the discrete topology. For each $f:X\rightarrow\mathbb{R}^{\mathcal{P}}$ denote $\hat{f}(P,x):=\pi_{P}\circ f(x)$ for all $(P,x)\in\mathcal{P}\times X$. Similarly, for each $g:\mathcal{P}\times X\rightarrow\mathbb{R}$ denote $\check{g}(x)(P):=g(P,x)$. Given $A\subseteq(\mathbb{R}^{\mathcal{P}})^{X}$, we denote $\hat{A}$ as the set of all $\hat{f}$ such that $f\in A$. Note that the map $\left(\mathbb{R}^{\mathcal{P}}\right)^{X}\to\mathbb{R}^{\mathcal{P}\times X}$ given by $f\mapsto\hat{f}$ is a homeomorphism and its inverse is given by $g\mapsto\check{g}$.
Lemma 1.11.
Let $X$ be a Polish space and $\mathcal{P}$ be countable. Then, $f\in B_{1}(X,\mathbb{R}^{\mathcal{P}})$ if and only if $\hat{f}\in B_{1}(\mathcal{P}\times X)$.
Proof.
($\Rightarrow$) By Lemma 1.10, given an open set of reals $U$, we have $f^{-1}[\pi_{P}^{-1}[U]]$ is $F_{\sigma}$ for every $P\in\mathcal{P}$. Given that $\mathcal{P}$ is a discrete countable space, we observe that
$\hat{f}^{-1}[U]=\bigcup_{P\in\mathcal{P}}\left(\{P\}\times f^{-1}[\pi_{P}^{-1}[U]]\right)$
is an $F_{\sigma}$ as well.
($\Leftarrow$) By lemma 1.10 it suffices to show that $\pi_{P}\circ f\in B_{1}(X)$ for all $P\in\mathcal{P}$. Fix an open $U\subseteq\mathbb{R}$. Write $\hat{f}^{-1}[U]=\bigcup_{n\in\mathbb{N}}F_{n}$ where $F_{n}$ is closed in $\mathcal{P}\times X$. Then,
$(\pi_{P}\circ f)^{-1}[U]=\bigcup_{n\in\mathbb{N}}\{x\in X:(P,x)\in F_{n}\}$
which is $F_{\sigma}$. ∎
@@
Assume that $\mathcal{P}$ is countable, $X$ is a Polish space, and $A\subseteq C_{p}(X,\mathbb{R}^{\mathcal{P}})$ is such that $\pi_{P}\circ A$ is pointwise bounded for all $P\in\mathcal{P}$. The following are equivalent for every compact $K\subseteq X$:
@@
    $\overline{A|_{K}}\subseteq B_{1}(K,\mathbb{R}^{\mathcal{P}})$.
@@
    $\pi_{P}\circ A|_{K}$ satisfies the NIP for every $P\in\mathcal{P}$.
@@
[…]
$(\mathcal{P}\times K)\cap\bigcap_{n\in I}\hat{f_{n}}^{-1}(-\infty,a]\cap\bigcap_{n\notin I}\hat{f_{n}}^{-1}[b,\infty)=\emptyset$
Hence,
$K\cap\bigcap_{n\in I}(\pi_{P}\circ f_{n})^{-1}(-\infty,a]\cap\bigcap_{n\notin I}(\pi_{P}\circ f_{n})^{-1}[b,\infty)=\emptyset$
[…]
$K\cap\bigcap_{n\in E}(\pi_{P}\circ f_{n})^{-1}(-\infty,a]\cap\bigcap_{n\in F}(\pi_{P}\circ f_{n})^{-1}[b,\infty)=\emptyset$
Thus, $\pi_{P}\circ A|_{L}$ satisfies the NIP.
[…]
@@
2. Compositional computation structures: A structural approach
to floating-point computation
In this section, we connect function spaces with floating point computation. We start by summarizing some basic concepts from [alva2024approximability].
A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates. For a state $v\in L$, the type of $v$ is defined as the indexed family
@@ 2. Compositional computation structures: A structural approach
Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable. We think of each state $v\in L$ as being uniquely characterized by its type $\operatorname{tp}(v)$. Thus, in practice, we identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$. A typical case will be when $L=\mathbb{R}^{\mathbb{N}}$ or $L=\mathbb{R}^{n}$ for some positive integer $n$ and there is a predicate $P_{i}(v)=v_{i}$ for each of the coordinates $v_{i}$ of $v$. We regard the space of types as a topological space, endowed with the topology of pointwise convergence inherited from $\mathbb{R}^{\mathcal{P}}$. In particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
@@
Given a computation states structure $(L,\mathcal{P})$, any element of $\mathbb{R}^{\mathcal{P}}$ in the image of $L$ under the map $v\mapsto\operatorname{tp}(v)$ will be called a realized type. The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of types of $(L,\mathcal{P})$, denoted $\mathcal{L}$. Elements of $\mathcal{L}\setminus L$ will be called unrealized types.
In traditional, compact-valued, model theory, the space of types of a structure is viewed as a sort of compactification of the structure, and the compactness of type spaces plays a central role. However, here we are dealing with real-valued structures, and the space $\mathcal{L}$ defined above is not necessarily compact. To bypass this obstacle, we follow the idea introduced in [alva2024approximability] of covering $\mathcal{L}$ by “thin” compact subspaces that we call shards. The formal definition of shard is next.
@@
A sizer is a tuple $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, we define the $r_{\bullet}$-shard as:
$L[r_{\bullet}]=L\cap\prod_{P\in\mathcal{P}}[-r_{P},r_{P}].$
For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$. We define $\mathcal{L}_{\text{sh}}$, as the union of all type-shards.

### v2

@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
    AA has BFT-NIP;
  3. (iii)
    AA has countable NIP;
  4. (iv)
    AA has finitary NIP.
[…]
Proof.
Corollary of Theorem 1.5 and Proposition 1.7. ∎
Theorem 1.8 may be stated as the following dichotomy (under the assumptions): either $A$ is relatively compact in $B_{1}(X)$, or $A$ has IP (in either sense).
The Independence Property was first isolated by Saharon Shelah in model theory as a dividing line between theories whose models are “tame” (corresponding to NIP) and theories whose models are “wild” (corresponding to IP). See [SHE71, Definition 4.1], [SHE90]. We will discuss this dividing line in more detail in the next section.
@@ 1.2. NIP as a universal dividing line between polynomial and exponential complexity
The particular case of the BFT dichotomy (Theorem 1.5) when $A$ consists of $\{0,1\}$-valued (i.e., $\{\text{Yes},\text{No}\}$-valued) strings was discovered independently, around 1971-1972 in many foundational contexts related to polynomial (“tame”) vs exponential (“wild”) complexity: In model theory, by Saharon Shelah [SHE71], [SHE90], in combinatorics, by Norbert Sauer [SAU72], and Shelah [SHE72], [SHE90], and in statistical learning, by Vladimir Vapnik and Alexey Chervonenkis [VC71], [VC74].
@@
In Shelah’s classification program [SHE90], theories without the independence property (called NIP theories, or dependent theories) have a well-behaved, “tame” structure; the number of types over a set of parameters of size $\kappa$ of such a theory is of polynomially or similar “slow” growth on $\kappa$.
In contrast, theories with the Independence Property (called IP theories) are considered “intractable” or “wild”. A theory with the Independence Property produces the maximum possible number of types over a set of parameters; for a set of parameters of cardinality $\kappa$, the theory will have $2^{2^{\kappa}}$-many distinct types.
@@
Sauer [SAU72] and Shelah [SHE72] proved the following independently: Let $\mathscr{F}$ be a family of subsets of some set $S$. Either: for every $n\in\mathbb{N}$ there is a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A:i\in\mathbb{N}\}|=2^{n}$ ($\mathscr{F}$ has “exponential complexity”); or: there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
$|\{S_{i}\cap A:i\in\mathbb{N}\}|\leq\sum_{i=0}^{N-1}\binom{|A|}{i}$.
($\mathscr{F}$ has “polynomial complexity”). Clearly, any family $\mathscr{F}$ of subsets of a finite set $S$ has polynomial complexity. The “polynomial” name is justified: indeed, for fixed $N>0$, as a function of the size $|A|=m>0$, we have
$\sum_{i=0}^{N-1}\binom{m}{i}\leq\sum_{i=0}^{N-1}\frac{m^{i}}{i!}\leq\left(\sum_{i=0}^{N-1}\frac{1}{i!}\right)\cdot m^{N-1}\cdot<e\cdot m^{N-1}=O\left(m^{N}\right).$
(More precisely, the order of magnitude is $O(m^{N-1})$: polynomial in $m$ for $N$ fixed.)
@@
Readers familiar with statistical learning may recognize the Sauer-Shelah lemma as the dichotomy discovered and proved slightly earlier (1971) by Vapnik and Chervonenkis [VC71, VC74] to address uniform convergence in statistics. The least integer $N$ given by the preceding paragraph, when it exists, is called the VC-dimension of $\mathscr{F}$; it is a core concept in machine learning. If such an integer $N$ does not exist, we say that the VC-dimension of $\mathscr{F}$ is infinite. The lemma provides upper bounds on the number of data points (sample size) needed to learn a concept class of known VC dimension $d$ up to a given admissible error in the statistical sense. The Fundamental Theorem of Statistical Learning states that a hypothesis class is PAC-learnable (PAC stands for “Probably Approximately Correct”) if and only if its VC dimension is finite.
@@ 1.3. Rosenthal compacta
The universal classification implied by Theorem 1.5, as attested by the examples outlined in the preceding section, led to the following definition (by Gilles Godefroy [GOD80]):
@@
A Rosenthal compactum is any topological space realized as a compact subset of the space $B_{1}(X)=B_{1}(X,\mathbb{R})$ (equipped with the topology of pointwise convergence) of all real functions of the first Baire class on some Polish space $X$.
A Rosenthal compactum $K$ is necessarily Hausdorff since it is a topological subspace of the Hausdorff product space $\mathbb{R}^{X}$.
Rosenthal compacta possess significant topological and dynamical tameness properties, and play an important role in functional analysis, measure theory, dynamical systems, descriptive set theory, and model theory. In this paper, we use them to study deep computations. For this, we shall first focus on countable languages, which is the theme of the next subsection.
@@
Fix an arbitrary (at most) countable set $\mathcal{P}$ whose elements $P\in\mathcal{P}$ will be called predicate symbols or formal predicates. Our present goal is to characterize relatively compact subsets of $B_{1}(X,\mathbb{R}^{\mathcal{P}})$, where $X$ is always assumed to be a perfectly normal space (often a Polish space).
The set $\mathcal{P}$ shall be considered discrete whenever regarded as a topological space. Since $C_{p}(X,\mathbb{R}^{\mathcal{P}})\subseteq B_{1}(X,\mathbb{R}^{\mathcal{P}})\subseteq(\mathbb{R}^{\mathcal{P}})^{X}$, the “ambient” space $(\mathbb{R}^{\mathcal{P}})^{X}$ is quite relevant. The product $X\times\mathcal{P}$ will be regarded as either a pointset, or as a topological product depending on context. We have natural homeomorphic identifications
$(\mathbb{R}^{\mathcal{P}})^{X}\cong\mathbb{R}^{X\times\mathcal{P}}\cong(\mathbb{R}^{X})^{\mathcal{P}},$
given by
$\mathbb{R}^{X\times\mathcal{P}}\to(\mathbb{R}^{\mathcal{P}})^{X}:\varphi$ $\mapsto{\hat{}}\varphi$
$\mathbb{R}^{X\times\mathcal{P}}\to(\mathbb{R}^{X})^{\mathcal{P}}:\varphi$ $\mapsto\varphi{\,\hat{}}$,
where
$\hat{}\varphi(x)\coloneqq\varphi(x,\cdot)\in\mathbb{R}^{\mathcal{P}},$ $\varphi\,\hat{}(P)\coloneqq\varphi(\cdot,P)\in\mathbb{R}^{X}.$
Such identifications view $X$ and $\mathcal{P}$ as mere pointsets (the topology of $X$ in particular plays no role).
For $x\in X$, define the “left projection” map
$\lambda_{x}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{\mathcal{P}}:\varphi\mapsto\lambda_{x}(\varphi)\coloneqq\varphi(x,\cdot);$
for $P\in\mathcal{P}$, the “right projection” map
$\rho_{P}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{X}:\varphi\mapsto\varphi(\cdot,P).$
For fixed $x\in X$ and $P\in\mathcal{P}$, we also have canonical projection maps
$\pi_{x}:\mathbb{R}^{X}\to\mathbb{R}:f\mapsto f(x),$ $\pi_{P}:\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:f\mapsto f(P).$
[…]
The Proposition below reduces the study of $\mathbb{R}^{\mathcal{P}}$-valued continuous or Baire-1 functions on $X$ to the special case of real-valued ones.
Proposition 1.10.
The identification $(\mathbb{R}^{\mathcal{P}})^{X}\cong\mathbb{R}^{X\times\mathcal{P}}\cong(\mathbb{R}^{X})^{\mathcal{P}}$ induces identifications
$C_{p}(X,\mathbb{R}^{\mathcal{P}})\cong C_{p}(X\times\mathcal{P})\cong C_{p}(X)^{\mathcal{P}},$ $B_{1}(X,\mathbb{R}^{\mathcal{P}})\cong B_{1}(X\times\mathcal{P})\cong B_{1}(X)^{\mathcal{P}}.$
(The cardinality of $\mathcal{P}$ plays no role.)
@@
The identification of $C_{p}$-spaces follows trivially from the definition of topological product and the fact that $\mathcal{P}$ is discrete: a continuous map $X\to\mathbb{R}^{\mathcal{P}}$ is precisely a $\mathcal{P}$-tuple of continuous functions $X\to\mathbb{R}$, and these correspond to continuous functions $X\times\mathcal{P}\to\mathbb{R}$. The identification of Baire-1 spaces follows immediately, since it is defined in terms of the purely topological notion of limit (in the ambient space) of sequences of continuous functions. ∎
@@
Assume that $\mathcal{P}$ is countable, $X$ is a Polish space, and $A\subseteq C_{p}(X,\mathbb{R}^{\mathcal{P}})$ is pointwise bounded in the sense that $\pi_{P}\circ A$ ($\subseteq C_{p}(X)$) is pointwise bounded for every $P\in\mathcal{P}$. The following are equivalent for every compact $K\subseteq X$:
@@
    $A|_{K}$ is relatively compact in $B_{1}(K,\mathbb{R}^{\mathcal{P}})$;
@@
    $\pi_{P}\circ A|_{K}$ has NIP for every $P\in\mathcal{P}$.
@@
Compact subsets $K\subseteq X$ are closed, hence also Polish. Therefore, the asserted equivalence follows from Theorems 1.5 and 1.7. ∎
@@
2. Compositional Computation Structures: A structural approach to floating-point computations
In this section, we connect function spaces with floating point computations. We start by summarizing some basic concepts from [ADI+24].
A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates. For a state $v\in L$, the type of a state $v$ is the indexed family
@@ 2. Compositional Computation Structures: A structural approach to floating-point computations
Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable. Each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$. Important space states are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$. We regard the space of types as a topological space, endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$. Via the identification $v\mapsto\operatorname{tp}(v)$, the states space $L$ is correspondingly topologized; in particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
@@
Given a computation states structure $(L,\mathcal{P})$, any element of $\mathbb{R}^{\mathcal{P}}$ in the image of $L$ under the map $v\mapsto\operatorname{tp}(v)$ will be called a realized (state) type. The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of types of $(L,\mathcal{P})$, denoted $\mathcal{L}$; elements $\xi\in\mathcal{L}$ are called state types. Elements of $\mathcal{L}\setminus L$ will be called unrealized types.
Intuitively, state types capture a notion of “ultrastate”.
The traditional setting of discrete first-order logic may be regarded as the special case in which predicates are binary (such predicates may take only the values 0 and $1$, say). In such a traditional setting, the space of types of a structure is necessarily compact; for some purposes, it serves as a sort of compactification of the structure itself. Compactness in such setting is critical. However, predicates of a computational states structure take values in the non-compact space $\mathbb{R}$, so the corresponding type space $\mathcal{L}$ need not be compact. To bypass this obstacle, we follow the idea, introduced in [ADI+24], of covering $\mathcal{L}$ by “thin” compact subspaces called shards.
@@
A sizer is a tuple $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, let $\mathbb{R}^{[r_{\bullet}]}=\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of a states space $L$ be
$L[r_{\bullet}]=L\cap\mathbb{R}^{[r_{\bullet}]}.$
For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$ (a closed, hence compact subset of $\mathbb{R}^{[r_{\bullet}]}$).
Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards (as the sizer $r_{\bullet}$ varies).
In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$ may be a proper subset. Equality holds in the important special case when $\mathcal{P}$ is countable (see *** below).